# LoTA-QAF: Lossless Ternary Adaptation for Quantization-Aware Fine-Tuning

### v1

@@ 2 Related Work
Figure 2: Comparison of prior adaptation methods with LoTA-QAF.
@@ 3.1 Preliminaries
[…]
@@ 3.2 Lossless Ternary Adaptation
Notably, the product of the ternary adapters $\mathbf{A}_{\text{T}}$ and $\mathbf{B}_{\text{T}}$ forms the auxiliary matrix $\Delta\mathbf{W}=\mathbf{A}_{\text{T}}\mathbf{B}_{\text{T}}$. Its elements $\Delta\mathbf{W}_{ij}$ are integers within the range $[-r,r]$, where $i\in\{1,\dots,D_{\text{in}}\}$ and $j\in\{1,\dots,D_{\text{out}}\}$. The auxiliary matrix $\Delta\mathbf{W}$ is mapped to the ternary matrix $\mathbf{\hat{W}}\in\{-1,0,1\}^{D_{\text{in}}\times D_{\text{out}}}$ using a threshold $\omega\in(0,r)$, as follows:
@@
$\mu$ $=\frac{\sum_{i=1}^{D_{\text{in}}}\sum_{j=1}^{D_{\text{out}}}\mathbf{\tilde{W}}_{ij}}{rD_{\text{in}}D_{\text{out}}},$
@@
where the final weights $\mathbf{W^{\prime}_{\text{int}}}$ belong to the set $\{0,1,\dots,2^{N}-1\}$. Therefore, the lossless ternary adaptation preserves low-bit computational efficiency and avoids the reintroduction of quantization loss at the adapter level.
Figure 3: Illustration of LoTA-QAF. The process is demonstrated using a small 4x4 matrix under 4-bit quantization, with adapter rank $r$ set to 3 and threshold $\omega$ to 1. Notably, quantized weights have boundary values (e.g., 0 and 15 for 4-bit), requiring overflow prevention.
@@ 3.3 Ternary Signed Gradient Descent
Inspired by the application of signed gradient descent (SignSGD) for quantization-related parameter updates in AutoRound (Cheng et al., 2023), we propose t-SignSGD for optimizing ternary adapters. AutoRound successfully employs SignSGD to navigate parameter updates within strictly bounded spaces. Indeed, Balles et al. (2020); Safaryan and Richtárik (2021); Li et al. (2023a, 2024) demonstrate that signSGD is beneficially applied to updates within discrete and constrained domains. This characteristic is highly relevant to our ternary adapters, which are constrained to the discrete values $\{-1,0,1\}$. Specifically, our proposed t-SignSGD performs updates on the ternary adapters (e.g., $\mathbf{A}_{\text{T}}$). Let $g_{t}=\nabla_{\mathbf{A}_{\text{T}}}\mathcal{L}|_{\mathbf{A}_{\text{T},t}}$ denote the gradient at iteration $t$. To focus updates on prominent gradients, we employ a fixed minimum gradient threshold $\tau$ (e.g., $1\times 10^{-9}$) and a dynamic percentile-based threshold $\sigma_{t}$. The threshold $\sigma_{t}$ is determined at each iteration $t$ based on a specific percentile of the gradient magnitudes, selecting a certain proportion of the largest gradients for update. This proportion of selected gradients is initially the top 5% and then linearly decays to 0.01% during training. The t-SignSGD update is performed as follows:
@@
where the indicator function $\mathbb{I}_{(\cdot)}$ equals 1 if the condition ($|g_{t}|>\max(\tau,\sigma_{t})$) is satisfied, and 0 otherwise. The $\text{clip}(\cdot,-1,1)$ function ensures that the updated adapter weights $\mathbf{A}_{\text{T},t+1}$ are strictly constrained to ternary set $\{-1,0,1\}$.
Notably, Equation 6 does not use a learning rate to scale the sign of the gradient. Instead, if a ternary adapter weight $\mathbf{A}_{\text{T}}$ is selected for an update (i.e., its gradient magnitude $|g_{t}|$ exceeds the threshold $\max(\tau,\sigma_{t})$), its value is adjusted by $-\text{sign}(g_{t})$. Therefore, the threshold $\sigma_{t}$ play a crucial role in determining the selectivity of updates. By targeting only the top percentile of gradient magnitudes (e.g., the top 5%), $\sigma_{t}$ acts as an adaptive mechanism that focuses updates on the most salient ternary adapter weights at each iteration $t$. The t-SignSGD is viable for ternary adapters due to two primary factors. First, the discrete and strictly bounded nature of the ternary values $\{-1,0,1\}$ is well-suited for sign-based updates. Secondly, updates are not performed directly on the quantized weights. Instead, these updates influence the quantized weights through the ternary adapters. Further analysis of parameters and convergence is provided in Section 4.3.
@@ 4.1 Experimental setup
We benchmark LoTA-QAF against two state-of-the-art methods: LoRA (Hu et al., 2022) and QA-LoRA (Xu et al., 2023). For fine-tuning quantized models to recover performance and align them with specific tasks, LoRA remains a widely adopted method (applied to GPTQ-quantized models, it can be considered a QLoRA (Dettmers et al., 2023) variant). Although LoRA introduces 16-bit adapters, resulting in reduced inference efficiency, the 16-bit precision of its adapters also enables more precise adjustment of the quantized weights. Moreover, as the closest work to ours, QA-LoRA maintains low-bit inference efficiency by lossless merging the adapters into quantized weights.
@@ 4.3 Further Analysis of LoTA-QAF
Second, the dynamic percentile-based threshold $\sigma_{t}$ (where $t$ is the iteration step) acts as a learning rate in t-SignSGD, as explained in Equation 6. We conduct parameter experiments by selecting initial $\sigma_{t}$ values corresponding to the top $\{9.5\%,8.0\%,6.5\%,5.0\%,3.5\%,2.0\%\}$ of gradient magnitudes. For the 8B model, an initial $\sigma_{t}$ of 5.0% (representing the top 5.0% of gradients) consistently yields favorable results. Conversely, selecting smaller initial $\sigma_{t}$ values leads to a noticeable decline in performance for both 3-bit and 2-bit quantization. This observation can be attributed to the nature of task-specific fine-tuning, which necessitates more substantial adjustments to model parameters compared to recovery performance. This aligns with our general learning rate strategy, where the learning rate for recovery performance is set lower than that for task-specific fine-tuning. Further hyper-parameter analysis is provided in the Appendix B.
@@
Regarding convergence, we evaluate LoRA and LoTA on Llama 3.1 8B for SQL generation with 4, 3 and 2 bit-widths quantization. LoRA achieves best convergence across all three bit-widths. Notably, at 2-bit, LoRA reaches a convergence loss of $0.132$, whereas LoTA achieves $0.375$, showing LoRA’s 16-bit adapter stability. However, LoTA, optimized with t-SignSGD, also converges. For 4-bit and 3-bit quantization, its convergence loss differs from LoRA $<0.01$. Considering that LoTA can be losslessly merged into the quantized weights, whereas LoRA cannot, this makes LoTA a highly competitive method, especially when deployment efficiency is a priority.
@@ Appendix A Implementation of Ternary Adapters
Notably, the implementation of ternary adapters faces challenges primarily due to framework limitations. Methodologically, LoTA-QAF employs ternary adapters, which offer significant advantages in storage and computational efficiency over 16-bit adapters. Consequently, further engineering for production-level deployment is valuable and represents a promising avenue for future work (see Section E). Currently, the training costs are acceptable with an implementation based on PyTorch and accelerated by Triton. A detailed discussion on training efficiency is available in Section C.
@@ Appendix C Training Efficiency Analysis
Our simulated ‘ternary‘ adapter utilizes ‘bfloat16‘ for implementation due to the lack of native support for the ‘ternary‘ dtype within the PyTorch framework. Consequently, LoTA-QAF does not exhibit a training efficiency advantage over LoRA. Specifically, LoTA requires 14.1%-25.4% more training time compared to LoRA. This increased training time is primarily attributed to the implementation of Eq. (4) and the process of aligning with the dequantization process of GPTQModel. Despite this higher training cost, LoTA’s ternary adapters can be losslessly merged into quantized weights, a capability that LoRA does not offer. In terms of memory, LoTA incurs an additional 2.6%-6.3% overhead compared to LoRA. This overhead derives from the boundary checks detailed in Section A, though the overall cost remains modest. Section E discusses feasible directions for further enhancing LoTA’s training efficiency.
Appendix D Limitations
@@
Appendix E Future Work
@@
Our exploration of t-SignSGD has been limited. Further research could focus on incorporating momentum mechanisms, refining constraints on t-SignSGD, and applying optimization techniques (e.g., cosine annealing, cyclical learning rates, or adaptive step-size strategies). In t-SignSGD, we replaced the learning rate with a dynamic percentile-based threshold, $\sigma_{t}$, to select the top-$x$of gradient signs for directly updating the ternary adapter weights within the set $\{-1,0,1\}$. Under this mechanism, a more detailed analysis of gradient distribution characteristics could be pursued to guide the selection of ternary adapter weights for updates, potentially offering improved convergence properties.

### v2

@@ 2 Related Work
Figure 2: Comparison of prior adaptation methods with LoTA-QAF. A merge can result in efficiency loss, where model weights are de-quantized to preserve the adapter’s precision but sacrifice inference efficiency, or accuracy loss, where the adapter’s precision is sacrificed to maintain low-bit inference. A lossless merge avoids this trade-off because adapter weights are utilized identically during both the training and inference (after merge), ensuring the adapter’s learned effect is fully preserved.
@@ 3.1 Preliminaries
where the scaling factor $s=(\max(\mathbf{W})-\min(\mathbf{W}))/(2^{N}-1)$ and the zero factor $z=\min(\mathbf{W})$. The symbol $\lfloor\cdot\rceil$ denotes rounding to the nearest integer. All elements in the integer matrix $\mathbf{W_{\text{int}}}$ belong to the set $\{0,1,\dots,2^{N}-1\}$.
@@ 3.2 Lossless Ternary Adaptation
Figure 3: Illustration of LoTA-QAF. The process is demonstrated using a small 4x4 matrix under 4-bit quantization, with adapter rank $r$ set to 3 and threshold $\omega$ to 1. Notably, quantized weights have boundary values (e.g., 0 and 15 for 4-bit, where 15 cannot be incremented and 0 cannot be decremented), requiring overflow prevention.
Notably, the product of the ternary adapters $\mathbf{A}_{\text{T}}$ and $\mathbf{B}_{\text{T}}$ forms the auxiliary matrix $\Delta\mathbf{W}=\mathbf{A}_{\text{T}}\mathbf{B}_{\text{T}}$. Its elements $\Delta\mathbf{W}_{ij}$ are integers within the range [r,r][-r,r], where $i\in\{1,\dots,D_{\text{in}}\}$ and $j\in\{1,\dots,D_{\text{out}}\}$. The auxiliary matrix $\Delta\mathbf{W}$ is mapped to the ternary matrix $\mathbf{\hat{W}}\in\{-1,0,1\}^{D_{\text{in}}\times D_{\text{out}}}$ using a threshold $\omega\in(0,r)$, as follows:
@@
$\mu$ $=\frac{\sum_{i=1}^{D_{\text{in}}}\sum_{j=1}^{D_{\text{out}}}\mathbf{\tilde{W}}_{ij}}{D_{\text{in}}D_{\text{out}}},$
@@
where the final weights $\mathbf{W^{\prime}_{\text{int}}}$ belong to the set $\{0,1,\dots,2^{N}-1\}$. Equivalently, the forward pass of the unmerged module is defined as $\mathbf{y}=(s\cdot\mathbf{W^{\prime}}_{\text{int}}+z^{\prime})^{T}\mathbf{x}$. Therefore, the lossless ternary adaptation preserves low-bit computational efficiency and avoids the reintroduction of quantization loss at the adapter level.
@@ 3.3 Ternary Signed Gradient Descent
Inspired by the application of signed gradient descent (SignSGD) for quantization-related parameter updates in AutoRound (Cheng et al., 2023), we propose t-SignSGD for optimizing ternary adapters. Indeed, Balles et al. (2020); Safaryan and Richtárik (2021); Li et al. (2023a, 2024) demonstrate that SignSGD is beneficially applied to updates within discrete and constrained domains. This characteristic is highly relevant to our ternary adapters, which are constrained to the discrete values $\{-1,0,1\}$. Specifically, our proposed t-SignSGD performs updates on the ternary adapters (e.g., $\mathbf{A}_{\text{T}}$). Let $g_{t}=\nabla_{\mathbf{A}_{\text{T}}}\mathcal{L}|_{\mathbf{A}_{\text{T},t}}$ denote the gradient at iteration $t$. To focus updates on prominent gradients, we employ a fixed minimum gradient threshold $\tau$ (e.g., $1\times 10^{-9}$) and a dynamic percentile-based threshold $\sigma_{t}$. The threshold $\sigma_{t}$ is determined at each iteration $t$ based on a specific percentile of the gradient magnitudes, selecting a certain proportion of the largest gradients for update. This proportion of selected gradients is initially the top 5% and then linearly decays to 0.01% during training. The t-SignSGD update is performed as follows:
@@
where the indicator function $\mathbb{I}_{(\cdot)}$ equals 1 if the condition ($|g_{t}|>\max(\tau,\sigma_{t})$) is satisfied, and 0 otherwise. The $\text{clip}(\cdot,-1,1)$ function ensures that the updated adapter weights $\mathbf{A}_{\text{T},t+1}$ are strictly constrained to ternary set $\{-1,0,1\}$. Notably, Eq. (6) does not use a learning rate to scale the sign of the gradient. Instead, if a ternary adapter weight $\mathbf{A}_{\text{T}}$ is selected for an update (i.e., its gradient magnitude $|g_{t}|$ exceeds the threshold $\max(\tau,\sigma_{t})$), its value is adjusted by $\text{sign}(g_{t})$. Therefore, the threshold $\sigma_{t}$ play a crucial role in determining the selectivity of updates. By targeting only the top percentile of gradient magnitudes (e.g., the top 5%), $\sigma_{t}$ acts as an adaptive mechanism that focuses updates on the most salient ternary adapter weights at each iteration $t$.
The convergence of t-SignSGD is primarily governed by its update rule, which is designed for the discrete and bounded nature of the ternary adapters. The key component is the indicator function $\mathbb{I}_{|g_{t}|>\max(\tau,\sigma_{t})}$ in Eq. (6), which acts as an adaptive and dynamic selection mechanism. It ensures stable convergence in two ways: i) Stability via noise filtering (preventing oscillation). In discrete optimization, small, noisy gradients can cause parameters to oscillate unstably between states (e.g., flipping between 0 and 1). Our thresholding mechanism filters out these low-magnitude gradients, ensuring that updates are only performed when the gradient signal is strong and reliable. This addresses the core challenge in sign-based methods related to low signal-to-noise ratios, effectively preventing unstable "jitter" and promoting a smoother convergence path. ii) Convergence via annealing-like dynamics. The decaying percentile threshold $\sigma_{t}$ introduces an annealing-like, coarse-to-fine search strategy. Early in training ($\sigma_{t}$ is high), the top percentile of gradients triggers updates. This focuses the optimization on the most impactful parameters, allowing the model to quickly perform "broad-stroke" adjustments and find a promising region in the vast solution space. Later in training ($\sigma_{t}$ is low), allowing for more fine-grained adjustments. This enables the model to refine its solution within the promising region, analogous to the exploitation phase in classic search algorithms. This dynamic balance between exploration and exploitation is a well-known strategy for improving convergence in complex landscapes. Further analysis of parameters and convergence is provided in Section 4.3.
@@ 4.1 Experimental setup
We benchmark LoTA-QAF against two state-of-the-art methods: LoRA (Hu et al., 2022) and QA-LoRA (Xu et al., 2023). For fine-tuning quantized models to recover performance and align them with specific tasks, LoRA remains a widely adopted method (applied to GPTQ-quantized models, it can be considered a QLoRA (Dettmers et al., 2023) variant). Although LoRA introduces 16-bit adapters, resulting in reduced inference efficiency, the 16-bit precision of its adapters also enables more precise adjustment of the quantized weights. Moreover, as the closest work to ours, QA-LoRA maintains low-bit inference efficiency by lossless merging the adapters into zero factors.
@@ 4.3 Further Analysis of LoTA-QAF
Second, the dynamic percentile-based threshold $\sigma_{t}$ (where $t$ is the iteration step) acts as a learning rate in t-SignSGD, as explained in Equation 6. We conduct parameter experiments by selecting initial $\sigma_{t}$ values corresponding to the top $\{9.5\%,8.0\%,6.5\%,5.0\%,3.5\%,2.0\%\}$ of gradient magnitudes. For the 8B model, an initial $\sigma_{t}$ of 5.0% (representing the top 5.0% of gradients) consistently yields favorable results. Conversely, selecting smaller initial $\sigma_{t}$ values leads to a noticeable decline in performance for both 3-bit and 2-bit quantization. This observation can be attributed to the nature of task-specific fine-tuning, which necessitates more substantial adjustments to model parameters compared to recovery performance. This aligns with our general learning rate strategy for LoRA and QA-LoRA, where the learning rate for recovery performance is set lower than that for task-specific fine-tuning. Further hyper-parameter analysis is provided in the Appendix B.
@@
Regarding convergence, we evaluate LoRA and LoTA on Llama 3.1 8B for SQL generation with 4, 3 and 2 bit-widths quantization. LoRA achieves best convergence across all three bit-widths. Notably, at 2-bit, LoRA reaches a convergence loss of $0.132$, whereas LoTA achieves $0.375$, showing LoRA’s 16-bit adapter stability. However, LoTA, optimized with t-SignSGD, also converges. For 4-bit and 3-bit quantization, its convergence loss differs from LoRA <0.01<0.01. Considering that LoTA can be losslessly merged into the quantized weights, whereas LoRA cannot, this makes LoTA a highly competitive method, especially when deployment efficiency is a priority.
@@ 5 Conclusion
6 Acknowledgments
The work of Long Shi was supported by the National Natural Science Foundation of China under Grant 62201475 and Sichuan Science and Technology Program under Grant 2024NSFSC1436. The work of Xuming Hu was supported by the National Natural Science Foundation of China (Grant No.62506318); Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); CAAI-Ant Group Research Fund; Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
@@ Appendix A Implementation of Ternary Adapters
Notably, the implementation of ternary adapters faces challenges primarily due to framework limitations. Methodologically, LoTA-QAF employs ternary adapters, which offer significant advantages in storage and computational efficiency over 16-bit adapters. Consequently, further engineering for production-level deployment is valuable and represents a promising avenue for future work (see Section F). Currently, the training costs are acceptable with an implementation based on PyTorch and accelerated by Triton. A detailed discussion on training efficiency is available in Section C.
@@ Appendix C Training Efficiency Analysis
Our simulated ‘ternary‘ adapter utilizes ‘bfloat16‘ for implementation due to the lack of native support for the ‘ternary‘ dtype within the PyTorch framework. Consequently, LoTA-QAF does not exhibit a training efficiency advantage over LoRA. Specifically, LoTA requires 14.1%-25.4% more training time compared to LoRA. This increased training time is primarily attributed to the implementation of Eq. (4) and the process of aligning with the dequantization process of GPTQModel. Despite this higher training cost, LoTA’s ternary adapters can be losslessly merged into quantized weights, a capability that LoRA does not offer. In terms of memory, LoTA incurs an additional 2.6%-6.3% overhead compared to LoRA. This overhead derives from the boundary checks detailed in Section A, though the overall cost remains modest. Section F discusses feasible directions for further enhancing LoTA’s training efficiency.
Appendix D Ternary Adapter Weight Changes under t-SignSGD Training
Table 2: Analysis of ternary adapter weight changes.
Metric Value Percentage Description
Total Parameters 167,772,160 100% Total adapter parameters across 32 layers. Adapters are applied to 7 key projections within the attention and MLP blocks of each layer.
Increased Parameters 28,952,566 17.26% Count of weights that changed in the positive direction.
Decreased Parameters 28,960,070 17.26% Count of weights that changed in the negative direction.
Unchanged Parameters 109,859,524 65.48% Count of weights that are unchanged.
To better understand the dynamics of our optimizer, we analyzed the parameter changes of the ternary adapters for Llama 3.1 8B after fine-tuning on the Alpaca dataset. Table 2 shows that the rate of change in the adapters is nearly identical in both positive and negative directions (17.26% vs. 17.27%). This suggests that the adaptation is not a simple unidirectional shift but a balanced, bidirectional fine-tuning process to fit the target task. Moreover, one-third of the parameters changed during adaptation, while two-thirds remained unchanged. This phenomenon is a result of the t-SignSGD design, specifically the adaptive threshold $\mathbb{I}_{|g_{t}|>\max(\tau,\sigma_{t})}$. This mechanism ensures our updates are based on a large gradient magnitude, which implies a high probability that the sign is correct. Conversely, it filters out updates associated with small gradient signals, which can be noisy or ill-scaled.
Appendix E Limitations
@@
Appendix F Future Work
@@
Our exploration of t-SignSGD has been limited. Further research could focus on incorporating momentum mechanisms, refining constraints on t-SignSGD, and applying optimization techniques (e.g., cosine annealing, cyclical learning rates, or adaptive step-size strategies). In t-SignSGD, we replaced the learning rate with a dynamic percentile-based threshold, $\sigma_{t}$, to select the top-xx% of gradient signs for directly updating the ternary adapter weights within the set $\{-1,0,1\}$. Under this mechanism, a more detailed analysis of gradient distribution characteristics could be pursued to guide the selection of ternary adapter weights for updates, potentially offering improved convergence properties.